OpenStudy (anonymous):

** NEED HELP WITH FRACTIONS, PLEASE!** As a part of a 3-day workout plan to prepare for a triathlon, an athlete ran 9 1/12 miles Monday, 12 1/7 miles Tuesday, and 15 1/21 Wednesday. How many miles did the athlete run in the 3 day period?

OpenStudy (anonymous):

you add them them together give me a minute and dang that guy runs a lot lol

OpenStudy (anonymous):

Find a common denominator

OpenStudy (anonymous):

yes im stuck lol

OpenStudy (anonymous):

\[9\frac{ 1 }{ 12 } + 12 \frac{ 1 }{ 7 } + 15 \frac{ 1 }{ 21 }\]

OpenStudy (anonymous):

You need to make all the numbers in the denominator the same when adding fractions

OpenStudy (anonymous):

What is one possible number that can be used as a common denominator?

OpenStudy (anonymous):

none nada

OpenStudy (anonymous):

cant find it

OpenStudy (anonymous):

Well that's why you need to find one EXAMPLE PROBLEM  What's a possible common denominator you could work with there?

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

its 36 23/84

OpenStudy (anonymous):

finally

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

thanks